OpenStudy (venomblast):

An urn contains 6 red balls and 3 blue balls. One ball is selected at random and is replaces by a ball of the= other color. A Second ball is then chosen. What is the conditional probability that the first all selected is red, given that the second ball was red?

OpenStudy (venomblast):

R_1 - First ball Red, R_2-second ball is red P(R_1|R_2)=P(R_1 AND R_2)/P(R_2) P(R_1 AND R_2) = 6/9 * 5/9 = 30/81=10/27 P(R_2) = 6/9 * 5/9 + 3/9 * 6/9 = 16/27 P(R_1|R_2)=(10/27)/(16/27) = 5/8. I got it wrong. Why?

OpenStudy (jonathan34):

well you need to use the red ball as as your second

OpenStudy (jonathan34):

if you know what i mean?

OpenStudy (venomblast):

but there only 6 red balls you dont care about the the 3 balls

OpenStudy (kropot72):

 \[\large P(R _{1}\cap R _{2})=\frac{6}{9}\times\frac{5}{9}=\frac{30}{81}\] \[\large P(B _{1}\cap R _{2})=\frac{3}{9}\times\frac{7}{9}=\frac{21}{81}\] \[\large P(second\ is\ red)=\frac{30}{81}+\frac{21}{81}=\frac{51}{81}\] \[\large P(first\ was\ red)=\frac{\frac{30}{81}}{\frac{51}{81}}=\frac{10}{17}\]

OpenStudy (venomblast):

You still haven't explain why is it 7/9

OpenStudy (kropot72):

If a blue ball is selected in the first draw, it is replaced by a red ball (the other color). This results in there being 2 blue balls and 7 red balls for the second draw. The probability of red in the second draw, in this case, is 7/9.

OpenStudy (venomblast):

There only 6 red balls.

OpenStudy (kropot72):

It should be obvious that the replacement balls are not part of the original 9 balls. So, if a blue ball is replaced by a red ball, that red ball is taken from a supply of extra balls. That is the reason for it being possible to have 7 red balls.

OpenStudy (venomblast):

I don't get it still

OpenStudy (kropot72):

What part don't you get?

OpenStudy (venomblast):

7/9

OpenStudy (kropot72):

Do you accept that there must be additional balls outside the urn to make the exchange of color possible? There is always a total of 9 balls inside the urn, however the ratio of red to blue balls changes after the first draw, when the exchange of color is made.

OpenStudy (kropot72):

If the first ball that is drawn is blue, then the blue ball is kept outside the urn and a red ball from the quantity initially outside the urn is put into the urn. This results in there being only 2 blue balls and 7 red balls inside the urn. Therefore when the second draw is made the probability of drawing a red ball is given by: \[\large P(red)=\frac{number\ of\ red\ balls}{total\ number\ of\ balls}=\frac{7}{9}\]

OpenStudy (venomblast):

so the blue ball turns to a red ball?

OpenStudy (kropot72):

Yes, if a blue ball is drawn first, the blue ball stays outside the urn and a red ball from a supply outside the urn is put into the urn.

OpenStudy (venomblast):

That makes sense. I thought you said that the color just turned but you really just adding the color back. So this questions is reallu asking with replacement.

OpenStudy (kropot72):

Yes, replacement with the other color.